### v2

@@ Acknowledgements.
- Cherrey et al. (2023) Cherrey, M., Bouché, N. F., Zabl, J., et al. 2023, MNRAS[arXiv:2312.01762]
@@ Appendix A Catalogs and data release
| SIGMA | Velocity dispersion (km s) of galaxies in the group, using the Gapper method |
@@
| LMS | Logarithm of stellar mass (in ) within a 3 arcsec aperture estimated from SED fitting with CIGALE, see Sect. 2.4 |
| DLMS | Uncertainty on the logarithm of stellar mass (in ) within 3 arcsec apertures |
| LSFR | Logarithm of SFR (in ) within a 3 arcsec aperture estimated from SED fitting with CIGALE, see Sect. 2.4 |
| DLSFR | Uncertainty on the logarithm of SFR (in ) within 3 arcsec apertures |
@@
| VMC_D | Galaxy density in log (Mpc) at the location of the galaxy inferred from the VMC method, see Sect. 4.3 |
@@
| VOR_D | Galaxy density in log (Mpc) at the location of the galaxy inferred from Voronoi tessellations, see Sect. 4.4 |
@@
| DV | Velocity (km s) in the frame of the group with respect to the group redshift, see Sect. 4.5 |

### v3

@@ Acknowledgements.
- Cherrey et al. (2024) Cherrey, M., Bouché, N. F., Zabl, J., et al. 2024, MNRAS, 528, 481
@@ Appendix A Catalogs and data release
| SIGMA | Velocity dispersion (km s-1) of galaxies in the group, using the Gapper method |
@@
| LMS | Logarithm of stellar mass (in ) within 3″ diameter apertures estimated from SED fitting with CIGALE, see Sect. 2.4 |
| DLMS | Uncertainty on the logarithm of stellar mass (in ) within 3″ diameter apertures |
| LSFR | Logarithm of SFR (in ) within 3″ diameter apertures estimated from SED fitting with CIGALE, see Sect. 2.4 |
| DLSFR | Uncertainty on the logarithm of SFR (in ) within 3″ diameter apertures |
@@
| VMC_D | Galaxy density in log (Mpc-2) at the location of the galaxy inferred from the VMC method, see Sect. 4.3 |
@@
| VOR_D | Galaxy density in log (Mpc-2) at the location of the galaxy inferred from Voronoi tessellations, see Sect. 4.4 |
@@
| DV | Velocity (km s-1) in the frame of the group with respect to the group redshift, see Sect. 4.5 |